The Village of Hawksbill issued $5,200,000 in 5 percent general obligation, tax-supported bonds on July 1, 2019, at 102. A fiPlease bold answer in explanation
Answer #1

Answer :-

A.

General Journal Debit Credit
2019
1 No Entry For Bond Sale– Recorded In Capital Projects Fund
Cash      $104,000
Other Financing Sources - Transfers In      $104,000
2 Cash    $130,000
Other Financing Sources - Transfers In    $130,000
Expenditures-Interest    $130,000
Cash $130,000
3 Other Financing Sources - Transfers In    $234,000
Expenditures-Interest    $130,000
Fund Balance      $104,000
2020
4 Cash    $286,000
Other Financing Sources - Transfers In    $286,000
Expenditures-Interest    $130,000
Expenditures-Principal    $260,000
Cash    $390,000
5 Cash      $123,500
Other Financing Sources - Transfers In      $123,500
Expenditures-Interest      $123,500
Cash      $123,500
6 Other Financing Sources - Transfers In    $409,500
Fund Balance      $104,000
Expenditures-Interest    $253,500
Expenditures-Principal    $260,000

Calculation

Entry #2 :

2019 Expenditures-Interest = =$5,200,000*5%*6/12 = $130,000

Entry #5 :

2020 Expenditures-Interest = ($5,200,000 - $260,000) * 5% * 6/12 = $123,500

B.

Village Of Hawksbill
Statement Of Revenues, Expenditures, And Changes In Fund Balance Year Ended December 31, 2019
Revenues:
Total Revenues 0
Debt Service Expenditures:
Interest $130,000
Principal 0
Excess (Deficiency) Of Revenues Over Expenditures -$130,000
Other Financing Sources:
Transfers In: $234,000
Total Change In Fund Balance $104,000
Fund Balance - January 1, 2019 0
Fund Balance - December 31, 2019 $104,000

Calculation

Excess (Deficiency) Of Revenues Over Expenditures = Total Revenues - Debt Service Expenditures = 0 - $130,000 = -$130,000

Total Change In Fund Balance = Excess (Deficiency) Of Revenues Over Expenditures + Transfers In = -$130,000 + $234,000 = $104,000

C.

Village Of Hawksbill
Statement Of Revenues, Expenditures, And Changes In Fund Balance Year Ended December 31, 2020
Revenues:
Total Revenues 0
Debt Service Expenditures:
Interest $253,500
Principal $260,000
Excess (Deficiency) Of Revenues Over Expenditures -$513,500
Other Financing Sources:
Transfers In: $409,500
Total Change In Fund Balance ($104,000)
Fund Balance - January 1, 2020 $104,000
Fund Balance - December 31, 2020 $0

Excess (Deficiency) Of Revenues Over Expenditures = Total Revenues - Debt Service Expenditures = 0 - ($253,500 +$260,000) = -$513,500

Total Change In Fund Balance = Excess (Deficiency) Of Revenues Over Expenditures + Transfers In = -$513,500 + $409,500 = -$104,000.